this problem has got me buzzing please help
To pickup a rare serum neede to treat a sick child major Evans flew from Tinden air force base to Capitol City and back. Going he went 420mph and returning 480mph.If the round trip took one and a half hours how long did it take to get there? and how long to return?and the distance city to city
go time-
return time-
distance city to city-

Answer by checkley71(8403)

420T=480(1.5-T) T BEING THE TIME FOR THE 420 MPH TRIP
SOLVING FOR T WE GET 420T=720-480T OR 900T=720 OR T=720/900 OR T=.8 HOURS
THE 420 MPH TRIP TOOK .8 HOURS. THE DISTANCE THEN IS 420X.8 OR 336 MILES
THE 480 MPH TRIP THE TOOK 1.5-.8 OR .7 HOURS. 480X.7=336 MILES.